The Securities and Exchange Commission has determined that the number of companies listed in NYSE declaring bankruptcy is approximately a Poisson distribution with a mean of 2.6 per month. Find the probability that more than 1 bankruptcy occur next month. Round your answer to four decimal places.

The Securities and Exchange Commission has determined that the number of companies listed in NYSE declaring bankruptcy is approximately a Poisson distribution with a mean of 2.6 per month. What is the expected value for the number of bankruptcies in the next 6 month period?

The Securities and Exchange Commission has determined that the number of companies listed in NYSE declaring bankruptcy is approximately a Poisson distribution with a mean of 2.6 per month. What is the standard deviation for the number of bankruptcies in the next 6 month period? Round your answer to 4 decimal places.

The Securities and Exchange Commission has determined that the number of companies listed in NYSE declaring bankruptcy is approximately a Poisson distribution with a mean of 2.6 per month. Find the probability that more than 5 bankruptcies occurs next two months. Round your answer to four decimal places.

During off hours, cars arrive at a tollbooth on the East-West toll road at an average rate of 0.5 cars per minute. The arrivals are distributed according to the Poisson distribution. What is the probability that during the next five minutes three cars will arrive? Round your answer to four decimal places.

During off hours, cars arrive at a tollbooth on the East-West toll road at an average rate of 0.5 cars per minute. The arrivals are distributed according to the Poisson distribution. What is the probability that at least one car will arrive in the next two minutes?
Answer #1
Answer #2

For the Poisson distribution, the probability of k successes with mean ? is given by:

P( k | ?) = {[e^(-?)] * (?^k)} / k!

Given ? = 2.6

27) Find the probability that no more than 1 bankruptcy occur next month.

P(k = 0 or 1 | 2.6) = P(0 | 2.6) + P(1 | 2.6)

= 0.074273578 + 0.193111303 ? 0.2674

The answer is E.
_______________

26) Find the probability that more than 1 bankruptcy occur next month.

P(k > 1 | 2.6) = 1 - P(0 | 2.6) - P(1 | 2.6)

= 1 - 0.2674 = 0.7326

Similarly others can be done.